Consider the following hypothetical problem. Potassium ions have a concentration of 115 x 10-3mol/liter inside the cell and 4.5 x 10-3 mol/liter outside the cell. The cell wall has a channel through which the potassium ions can travel and it is 25.0 nm in length and 4.00 nm in diameter. Take the diffusion coefficient of potassium to be 3.00 x 10-9 m2/s.

A)What is the difference in concentration, Δc across the cell wall in molecules/m3?

B)What is the cross-sectional area of the potassium channel?

C)What is the rate of molecule transfer of the potassium ions?
Answer #1

Consider the following hypothetical problem. Potassium ions have a concentration of 115 x 10-3mol/liter inside the cell and 4.5 x 10-3 mol/liter outside the cell. The cell wall has a channel through which the potassium ions can travel and it is 25.0 nm in length and 4.00 nm in diameter. Take the diffusion coefficient of potassium to be 3.00 x 10-9 m2/s.

A)What is the difference in concentration, Δc across the cell wall in molecules/m3?

B)What is the cross-sectional area of the potassium channel?

C)What is the rate of molecule transfer of the potassium ions?

A)

Ci = concentration inside the cell = 115 x 10-3 mol/liter = 115 x 10-3 mol/(10-3 m3) = 115 mol/m3

Co = concentration outside the cell = 4.5 x 10-3 mol/liter = 4.5 x 10-3 mol/(10-3 m3) = 4.5 mol/m3

Difference in concentration is given as

\DeltaC = Co - Ci = 4.5 - 115 = - 110.5 mol/m3

we know that 1 mol = 6.023 x 1023 molecules

\DeltaC = - 110.5 x 6.023 x 1023 molecules/m3

\DeltaC = - 6.66 x 1025 molecules/m3

B)

d = diameter of the channel = 4 nm = 4 x 10-9 m

Area of cross-section of the channel is given as

A = (0.25) \pid2

A = (0.25) (3.14) (4 x 10-9)2

A = 1.256 x 10-17 m2

c)

x = length = 25 x 10-9 m

D = diffusion constant = 3 x 10-9 m2/s

rate of transfer = D A \DeltaC /x

rate of transfer = (3 x 10-9) (1.256 x 10-17) (6.66 x 1025)/(25 x 10-9)

rate of transfer = 1.004 x 108